# Beyond Walking: A Large-Scale Image-Text Benchmark for Text-based Person Anomaly Search

### v1

@@ Abstract
Text-based person search aims to retrieve specific individuals across camera networks using natural language descriptions. However, current benchmarks often exhibit biases towards common actions like walking or standing, neglecting the critical need for identifying abnormal behaviors in real-world scenarios. To meet such demands, we propose a new task, text-based person anomaly search, locating pedestrians engaged in both routine or anomalous activities via text. To enable the training and evaluation of this new task, we construct a large-scale image-text Pedestrian Anomaly Behavior (PAB) benchmark, featuring a broad spectrum of actions, e.g., running, performing, playing soccer, and the corresponding anomalies, e.g., lying, being hit, and falling of the same identity. The training set of PAB comprises $1,013,605$ synthesized image-text pairs of both normalities and anomalies, while the test set includes $1,978$ real-world image-text pairs. To validate the potential of PAB, we introduce a cross-modal pose-aware framework, which integrates human pose patterns with identity-based hard negative pair sampling. Extensive experiments on the proposed benchmark show that synthetic training data facilitates the fine-grained behavior retrieval in the real-world test set, while the proposed pose-aware method further improves the recall@1 by $2.88\%$. We will release the dataset, code, and checkpoints to facilitate further research and ensure the reproducibility of our results.
@@
Datasets Data source #Images #Text #Avg Text Len. Resolution Anno. Type Action
CUHK-PEDES [24] Market [51], etc. 40,206 80,440 23.5 $246\times 90$ Appearance Normal
ICFG-PEDES [11] MSMT-17 [45] 54,522 54,522 37.2 $378\times 142$ Appearance Normal
RSTPReid [55] MSMT-17 [45] 20,505 41,010 25.8 $546\times 219$ Appearance Normal
MALS [48] Synthesis 1,510,330 1,510,330 27.0 $531\times 208$ Appearance & Normal
Attribute
[…]
Collection Action & Scene Anomaly
Table 1: Dataset Characteristic Comparison. Here we present a comparison between our proposed Pedestrian Anomaly Behavior (PAB) benchmark and other prevailing Text-based Re-ID datasets on the data quality and quantity. “Len.” and “Anno.” are the abbreviations of “Length” and “Annotation”.
Split Data #Images #Avg Text Normal vs.
Source Length Anomaly
Training Set Synthesis 1,013,605 52 2:3
Test Set Real-world 1,978 50 1:1
Table 2: Statistics of PAB. Training and test sets of PAB all include person image-text pairs with scene, action, and anomaly annotations. In the training set, we generate about $1M$ images, and the ratio of normal behavior images to anomaly ones is $2:3$. In the test set, we manually collect pedestrian images from videos, and the ratio of normal to anomaly is $1:1$.
@@ 1 Introduction
Text-based person search is a well-established task [24, 53] that involves retrieving specific individuals from a large-scale image database using natural language descriptions. The textual query is intuitive for user-interactive applications such as smart city, security, and personalized services, where image queries are not available or hard to access. Current benchmarks for text-based person search suffer from the bias towards common actions, leading to a lack of diversity in the types of behaviors. For instance, prevailing datasets like CUHK-PEDES [24], ICFG-PEDES [11] and RSTReid [55], primarily focus on pedestrian actions such as walking and standing, which do not capture the full spectrum of real-world activities (see Figure 1). Such bias limits the generalizability and applicability of models trained on these datasets, particularly in scenarios where detecting abnormal behaviors is crucial (see Table 1).
To address this limitation, we introduce a new task, i.e. text-based person anomaly search. This task extends the traditional scope of person search by requiring the identification of pedestrians involved in both routine and anomalous activities through natural language descriptions. For example, while existing methods could successfully locate a person walking, they often fail to identify the same person lying on the ground or being hit, which is critical for security events tracing and locating, emergency response, and many other security applications. The importance of text-based person anomaly search cannot be overstated. In real-world scenarios, abnormal behaviors such as falls, assaults, and accidents require immediate attention and intervention. However, traditional person search methods that focus on common actions are insufficient for these critical situations.
@@
To fill this gap, we propose the Pedestrian Anomaly Behavior (PAB) benchmark. The PAB dataset is designed to cover a broad spectrum of actions, including both routine activities and anomalies. Specifically, the dataset features a wide range of actions such as running, performing, and playing soccer, along with corresponding anomalies like lying, being hit, and falling. Each action and anomaly is associated with the same individual identity, ensuring that the dataset accurately reflects real-world scenarios. As shown in Table 2, the PAB benchmark consists of a large-scale training set with $1,013,605$ synthesized image-text pairs, representing both normal and anomalous behaviors. The test set includes $1,978$ real-world image-text pairs, providing a realistic evaluation environment. The inclusion of synthesized data ensures that the training set is diverse and comprehensive, while the real-world test set validates the model performance in practical settings. To validate the potential of the PAB benchmark, we introduce a Cross-Modal Pose-aware (CMP) framework that integrates human pose patterns with identity-based hard negative pair sampling. This framework leverages the rich pose information to better distinguish between normal and anomalous behaviors. Our extensive experiments show that synthetic training data significantly enhances the fine-grained behavior retrieval performance on the real-world test set. Furthermore, the proposed pose-aware method achieves a substantial improvement. Therefore, our primary contributions are:
  • We introduce the task of text-based person anomaly search, which aims to locate pedestrians engaged in both routine and anomalous activities using natural language descriptions. Given the lack of a dataset in this field, we construct the Pedestrian Anomaly Behavior (PAB) benchmark, featuring $1,013,605$ synthesized and $1,978$ real-world image-text pairs with a broad spectrum of actions and anomalies.
  • We propose a Cross-Modal Pose-aware (CMP) framework that integrates human pose patterns with identity-based hard negative pair sampling to enhance the discrimination between normal and anomalous behaviors. This framework leverages structural information from human poses to improve the understanding of pedestrian activities, leading to better retrieval performance.
@@
  • Extensive experiments on the PAB benchmark show that synthetic training data effectively facilitates fine-grained behavior retrieval in real-world test sets. Our pose-aware method further improves recall@1 by $+2.88\%$, highlighting its effectiveness in identifying anomalous behaviors.
@@ 2 Related Work
Text-based Person Retrieval. Text-based pedestrian retrieval is introduced by Li et al. [24], which is a pioneering work to incorporate textual queries into large-scale pedestrian retrieval tasks, breaking the limitations of image-based [3, 33, 52] or attribute-based queries [27, 18, 28]. Text-based pedestrian retrieval is more challenging than general image-based pedestrian retrieval [3, 33, 52] and general cross-modal retrieval tasks [47, 49, 12, 17] due to its cross-modal and fine-grained nature. The key of this task is to learn the alignment of the image and text [54, 9, 34, 25, 16]. Roughly, prevailing alignment strategies can be divided into cross-modal attention-based [24, 43, 39, 38] and cross-modal non-attention [11, 54, 42, 9] methods. With the development of vision-language pre-training models [35, 23], recent works [19, 39, 46, 22, 5] have improved the robustness of learned features by transferring knowledge from large-scale generic image-text pairs. Inspired by the large language and multimodal models, recent studies attempt to utilize the large model to facilitate this problem [48, 41]. Yang et al. [48] presents a large-scale synthetic image-text dataset named MALS for pre-training, and Tan et al. [41] generate a large-scale LUPerson-MLLM dataset with Multimodal Large Language Models (MLLMs) to study transferable text-to-image ReID problems. In this paper, we introduce Text-based Person Anomaly Search, focusing on fine-grained behavior retrieval.
@@
Person Anomaly Detection. Existing work on pedestrian anomaly detection typically refers to Video Anomaly Detection (VAD). The goal of VAD is to detect events in videos that deviate from normal patterns, and it has been studied under various settings: as a one-class classification problem [14, 15, 21, 36, 50] where training only involves normal data; as an unsupervised learning task [50] where anomalies exist in the training set but it is unknown which training videos contain them; and as a supervised or weakly supervised problem [50, 1, 40] where training labels indicate anomalous video frames or videos containing anomalies. Most of these works focus on the one-class classification problem. Consequently, many video datasets are limited in terms of the number of videos or the realism of the scenarios. Examples include the UCSD Ped1 and Ped2 [26], Avenue [31], Subway [2], ShanghaiTech Campus [32], UBnormal [1], and NWPU [6]. Although Sultani et al. [40] construct a realistic video dataset called UCF-Crime, which annotates $13$ pre-defined categories of video anomalies, it still falls short in handling complex real-world scenarios. Therefore, we propose a large-scale multi-modal Pedestrian Anomaly Behavior (PAB) dataset that contains large-scale diverse normal and anomalous image-text pairs, capable of addressing more complex multi-modal anomaly behavior retrieval.
Figure 2: Dataset Properties (left). Compared with existing datasets for person re-ID, attribute recognition, text-based person search, and anomaly action recognition, ours contain more detailed action and appearance descriptions for text-based anomaly search. Dataset Examples (right). Our training set is synthesized, while the test set is collected from real-world videos. We provide the pair-wise training samples in terms of normal action and anomaly action.
@@ 3.1 Real-world Test Data Collection
We build our test benchmark from real-world videos to align with the practical applications. The OOPS! video [13] is chosen for its real-world recording, varied actions, and diverse scenes.
@@
  • Subtle discrepancy between image pairs. To overcome this weakness, we use ResNet-50 [20] to extract features from the images and calculate the cosine similarity between normal and anomaly images, filtering out pairs with a similarity greater than $0.95$.
  • False anomaly images, i.e., behaviors in anomaly images are normal. For this problem, we conduct manual verification by three professionals with advanced education in computer science, retaining only those image pairs of certain normal and anomaly behaviors. Through the aforementioned steps, we obtain $989$ high-quality image pairs ($1,978$ pedestrian images), which are designated as the normal and anomaly image pairs for the test set.
Caption Generation and Quality Control. Each OOPS! video is annotated with two types of captions, one ($C_{n}$) for normal moment and another ($C_{a}$) for anomaly occurring. Directly adopting these captions as image captions does not work well, because most of the captions are short and without descriptions of appearance. It is hard to retrieve the corresponding pedestrians by such text queries. Drawing inspiration from the significant progress in the Multi-modal Large Language Model (MLLM) [4, 8], we automatically generate captions for each image, eliminating the need for costly manual annotation. Particularly, we choose Qwen2-VL [4] as Image Captioner. The specific Instruction is as follows: “Provide a simple description of the image content within 50 words, including the appearance, attire, and actions of the main figures in the picture. Do not imagine any contents that are not in the image. Do not describe the atmosphere of the image.” The captions generated by MLLM can generally provide accurate descriptions of the appearance and actions of people in the images, including detailed information. However, due to inherent limitations of the model, the captions may contain minor errors. Therefore, we conduct Human Quality Control on the test set. Specifically, we enlist three professionals with advanced degrees in computer science to manually correct the text data. This measure ensures that the texts accurately match the content of the images and that the captions for normal and anomaly image pairs precisely capture specific normal behaviors and anomaly behaviors.
@@ 3.2 Large-scale Training Data Generation
Pedestrian-focused Image Generation. To ensure the generated images are realistic and stylistically consistent with the test set, we use captions from the OOPS! dataset. The $C_{n}$ and $C_{a}$ captions are input into the Realistic Vision V4.0 model [37] to generate high-quality pedestrian images. It is worth noting that, before image generation, we filter the captions to ensure they describe human subjects. For each caption, we extract the subject and verify if it refers to a person, such as “people”, “he”, “man”, etc. This process yields $6,739$ $C_{n}$ and $6,979$ $C_{a}$ captions. Additionally, for cases where both $C_{n}$ and $C_{a}$ are retained, we concatenate them to form a new $C_{a+}$. The resulting $6,669$ $C_{a+}$ captions serve as prompts for generating pedestrian images with anomaly behaviors, enhancing diversity. Despite the ability of the Realistic Vision V4.0 model to produce photorealistic images, it occasionally generates images with unreasonable human structures. To mitigate this issue, we use OpenPose [7] for human key point detection and eliminate images with structural errors. After generating 75 images per caption and applying image filtering, we obtained 1,013,605 high-quality, diverse pedestrian images, with a normal to anomalous behavior image ratio of approximately 2:3.
Figure 3: (a) Overview of our Cross-Modal Pose-aware (CMP) framework, composed of (b) a pose-aware image encoder, a text encoder, and a cross encoder. It uses (c) Identity-based Hard Negative Mining to form challenging negative pairs during training, followed by feature extraction, contrastive learning, and processing by the cross encoder with ITM and MLM heads. The final step computes a cross-modal loss including ITM and MLM components.

### v2

@@ Abstract
Text-based person search aims to retrieve specific individuals across camera networks using natural language descriptions. However, current benchmarks often exhibit biases towards common actions like walking or standing, neglecting the critical need for identifying abnormal behaviors in real-world scenarios. To meet such demands, we propose a new task, text-based person anomaly search, locating pedestrians engaged in both routine or anomalous activities via text. To enable the training and evaluation of this new task, we construct a large-scale image-text Pedestrian Anomaly Behavior (PAB) benchmark, featuring a broad spectrum of actions, e.g., running, performing, playing soccer, and the corresponding anomalies, e.g., lying, being hit, and falling of the same identity. The training set of PAB comprises $1,013,605$ synthesized image-text pairs of both normalities and anomalies, while the test set includes $1,978$ real-world image-text pairs. To validate the potential of PAB, we introduce a cross-modal pose-aware framework, which integrates human pose patterns with identity-based hard negative pair sampling. Extensive experiments on the proposed benchmark show that synthetic training data facilitates the fine-grained behavior retrieval, and the proposed pose-aware method arrives at $84.93\%$ recall@1 accuracy, surpassing other competitive methods. The dataset, model, and code are available at https://github.com/Shuyu-XJTU/CMP.
@@ 1 Introduction
Datasets Modality Annotation Content #Frames #Texts #Action Types Anomaly:Normal Data source
CUHK-PEDES [24] Image, Text Frame-level Text Appearance 40,206 80,440 - 0 Collection
ICFG-PEDES [11] Image, Text Frame-level Text Appearance 54,522 54,522 - 0 Collection
RSTPReid [59] Image, Text Frame-level Text Appearance 20,505 41,010 - 0 Collection
UBnormal [1] Video Frame-level Tag Binary Label 236,902 - 22 Anomaly 2:3 Synthesis
ShanghaiTech [32] Video Frame-level Tag Binary Label 317,398 - 11 Anomaly 1:18 Collection
UCF-Crime [43] Video Video-level Tag Binary Label 13,741,393 - 13 Anomaly $\ll$1:1 Collection
UCA [53] Video, Text Video-level Text Action 13,741,393 23,542 13 Anomaly $\ll$1:1 Collection
PAB (Ours) Image, Text Frame-level Text Appearance, 1,015,583 1,015,583 1,600 Anomaly 3:2 Synthesis &
Action, Scene 1,000 Normal Collection
Table 1: Dataset Characteristic Comparison. We present a comprehensive comparison between our proposed Pedestrian Anomaly Behavior (PAB) benchmark and existing text-based pedestrian search and video anomaly detection datasets in terms of data quality and quantity. Our dataset addresses the long-tail distribution challenge by incorporating a higher number of anomaly cases, while offering frame-level annotations (appearance, action, and scene) and fine-grained action types.
@@
Text-based person search [24, 57] focuses on retrieving specific individuals from large-scale image databases using natural language descriptions. This capability is particularly valuable for user-interactive applications in smart cities, security systems, and personalized services, where image queries are often unavailable or impractical to obtain. However, current benchmarks for text-based person search suffer from significant limitations in behavioral diversity, primarily due to their bias towards common pedestrian actions. Existing datasets, including CUHK-PEDES [24], ICFG-PEDES [11], and RSTReid [59], predominantly feature routine activities such as walking and standing, failing to adequately represent the full spectrum of real-world behaviors (see Figure 1). This inherent bias significantly restricts the generalizability and practical applicability of models trained on these datasets, particularly in critical scenarios requiring anomaly behavior detection.
Moreover, current anomaly detection methodologies focus on identifying and classifying predefined events, while existing video datasets face three major limitations: (1) limited data volume, typically containing less than 500k frames [26, 31, 34], (2) coarse annotation granularity, often restricted to binary labels (anomaly vs. normal) [32, 43], and (3) behavior long-tail distribution, where anomaly key frames are significantly underrepresented compared to normal behaviors [2, 38, 53] (see Table 1). These limitations hinder the development of robust models capable of handling diverse real-world scenarios. While UBnormal [1] addresses multi-class annotations using simulated 3D environments, the gap between virtual scenes and real-world footage limits generalization in practical deployments. Moreover, real-world applications often demand the ability to search for specific behaviors rather than broad, predefined categories, further highlighting the need for more nuanced datasets.
@@
To address the limitations in both fields, we introduce a new task, i.e., text-based person anomaly search. (1) This task extends the traditional scope of text-based person search by requiring the identification of pedestrians involved in both routine and anomalous activities. While existing methods could successfully locate a person walking, they often fail to identify the same person lying on the ground or being hit. (2) Comparing with the conventional anomaly detection, the proposed task further explores a fine-grained anomaly understanding beyond the binary tags, which is critical for security events tracing and locating, emergency response, and many other security applications. To support this task, we propose the Pedestrian Anomaly Behavior (PAB) benchmark. As summarized in Table 1, the PAB benchmark consists of $1,015,583$ image-text pairs, each annotated with detailed textual descriptions of the target pedestrian’s appearance, actions, and surrounding scene. The dataset comprehensively spans 1,000 distinct normal action, such as running, performing, and playing soccer, along with 1,600 anomalies like lying, being hit, and falling. This broad coverage ensures diversity in both routine and rare pedestrian activities, while the intentionally elevated anomaly ratio (3:2) enhances its utility for training robust anomaly detection models. To validate the potential of the PAB benchmark, we introduce a Cross-Modal Pose-aware (CMP) framework that integrates human pose patterns with identity-based hard negative pair sampling. This framework leverages the rich pose information to distinguish between normal and anomalous behaviors. Extensive experiments show that synthetic training data facilitate the person anomaly search on the real-world test set. The proposed framework also achieves a substantial improvement. In a summary, our primary contributions are:
  • We pioneer the task of text-based person anomaly search that unifies detection of both routine and anomalous activities through natural language queries. To address the critical absence of fine-grained anomaly data, we construct the Pedestrian Anomaly Behavior (PAB) benchmark, featuring $1.01M$ image-text pairs spanning $1,6000$ anomaly and $1,000$ normal action types.
  • Departing from conventional identity-centric person search methods, we propose a Cross-Modal Pose-aware (CMP) framework that integrates text, image, and human pose patterns for representation learning. We also introduce identity-based hard negative mining by perturbing the action to distinguishing subtle behavioral differences.
@@
  • Extensive experiments show that (1) synthetic training data of the proposed PAB facilitates fine-grained behavior retrieval in the real-world test set; (2) our full CMP model achieves 84.93% recall@1 accuracy on PAB and 55.23% recall@1 accuracy on the out-of-distribution (OOD) UCC testing, outperforming multiple competitive approaches.
@@ 2 Related Work
Text-based Person Retrieval. Text-based pedestrian retrieval incorporates textual queries into large-scale pedestrian retrieval tasks, breaking the limitations of image-based [3, 33, 56] or attribute-based queries [27, 18, 28]. Text-based pedestrian retrieval is more challenging than general image-based pedestrian retrieval [3, 33, 56] and general cross-modal retrieval tasks [50, 52, 12, 17] due to its cross-modal and fine-grained nature. The key to this task is to learn the alignment of the image and text [58, 9, 35, 25, 16]. Roughly, prevailing alignment strategies can be divided into cross-modal attention-based [24, 46, 42, 41, 36] and cross-modal non-attention [11, 58, 45, 9] methods. With the development of vision-language pre-training models [37, 23], recent works [19, 42, 49, 22, 5] have improved the robustness of learned features by transferring knowledge from large-scale generic image-text pairs. Inspired by the large language and multimodal models, recent studies attempt to utilize the large model to facilitate this problem [51, 44]. Yang et al. [51] presents a large-scale synthetic image-text dataset MALS for pre-training, while Tan et al. [44] generate a large-scale LUPerson-MLLM dataset to study transferable text-to-image ReID problems. In this paper, we introduce text-based person anomaly search, focusing on fine-grained behavior retrieval.
@@
Person Anomaly Detection. Existing work on pedestrian anomaly detection typically refers to Video Anomaly Detection (VAD) [6]. The goal of VAD is to detect events in videos that deviate from normal patterns, and it has been studied under various settings: as a one-class classification problem [14, 15, 21, 39, 54] where training only involves normal data; as an unsupervised learning task [54] where anomalies exist in training set but it is unknown which training videos contain them; and as a supervised or weakly supervised problem [54, 1, 43] where training labels indicate anomalous video frames or videos containing anomalies. Most of these works focus on the one-class classification problem. Consequently, many video datasets are limited in terms of the number of videos or the realism of the scenarios. Examples include the UCSD Ped1 and Ped2 [26], Avenue [31], Subway [2], ShanghaiTech Campus [32], and UBnormal [1]. Sultani et al. [43] construct a realistic video dataset called UCF-Crime, which annotates $13$ pre-defined categories of video anomalies but still falls short in handling complex real-world scenarios. Despite efforts [53, 48] to annotate UCF-Crime videos with text, datasets such as UCA [53] remain constrained (see Table 1). Therefore, we propose a large-scale multi-modal Pedestrian Anomaly Behavior (PAB) dataset that contains large-scale diverse normal and anomalous image-text pairs, capable of addressing more complex multi-modal anomaly behavior retrieval.
@@ 3 Person Anomaly Benchmark
Figure 2: Dataset Properties (left). Compared with existing datasets for person re-ID, attribute recognition, text-based person search, and anomaly action recognition, ours contain more detailed action and appearance descriptions for text-based anomaly search. Dataset Examples (right). Our training set is synthesized, while the test set is collected from real-world videos. We provide similar training samples in terms of normal and anomaly action, facilitating effective comparative analysis during model training.
@@ 3.1 Real-world Test Data Collection
To establish a practical evaluation benchmark, we construct our test set using real-world videos from the OOPS! dataset [13]. Note that this serves as a preliminary raw collection. Next we will illustrate how to perform rigorous filtering and annotation refinement to ensure data quality and task alignment.
@@
  • Duplicated images or images with subtle discrepancy. To overcome such noise, we apply ResNet-50 [20] to extract features from the images and calculate the cosine similarity between normal and anomaly images, filtering out pairs with a similarity greater than $0.95$.
  • False-positive anomaly images, i.e., behaviors in anomaly images are normal. For this problem, we conduct manual verification by three professionals with advanced education in computer science, retaining only those image pairs of certain normal and anomaly behaviors. Through the aforementioned steps, we obtain $989$ high-quality image pairs ($1,978$ pedestrian images), which are designated as 1:1 normal and anomaly image pairs for the test set.
Caption Generation and Quality Control. Each OOPS! video is annotated with two types of captions, one ($C_{n}$) for normal moment and another ($C_{a}$) for anomaly occurring. Directly adopting these captions as image captions does not work well, because most of the captions are short and without detailed descriptions of appearance. Drawing inspiration from the significant progress in the Multi-modal Large Language Model (MLLM) [4, 8], we automatically generate captions for each image, eliminating the need for costly manual annotation. Particularly, we choose Qwen2-VL [4] as Image Captioner. The specific Instruction is as follows: “Provide a simple description of the image content within 50 words, including the appearance, attire, and actions of the main figures in the picture. Do not imagine any contents that are not in the image. Do not describe the atmosphere of the image.” The captions generated by MLLM can generally provide accurate descriptions of the appearance and actions of people in the images, including detailed information. However, due to inherent limitations of the model, the captions may contain minor errors. Therefore, we conduct Human Quality Control on the test set. Specifically, we enlist three professionals with advanced degrees in computer science to manually correct the text data. This measure ensures that the texts accurately match the image content especially for specific normal and anomaly behaviors.
@@ 3.2 Large-scale Training Data Generation
Pedestrian-focused Image Generation. To ensure the generated images are realistic and stylistically consistent with the test set, we use captions from the OOPS! dataset. The $C_{n}$ and $C_{a}$ captions are input into the Realistic Vision V4.0 model [40] to generate high-quality pedestrian images. It is worth noting that, before image generation, we filter the captions to ensure they describe human subjects. For each caption, we extract the subject and verify if it refers to a person, such as “people”, “he”, “man”, etc. This process yields $6,739$ $C_{n}$ and $6,979$ $C_{a}$ captions. Additionally, for cases where both $C_{n}$ and $C_{a}$ are retained, we concatenate them to form a new $C_{a+}$. The resulting $6,669$ $C_{a+}$ captions serve as prompts for generating pedestrian images with anomaly behaviors, enhancing diversity. Despite the ability of the Realistic Vision V4.0 model to produce photorealistic images, it occasionally generates images with unreasonable human structures. To mitigate this issue, we use OpenPose [7] for human key point detection and eliminate images with structural errors. After generating 75 images per caption and applying image filtering, we obtain 1,013,605 high-quality, diverse pedestrian training images, with an anomalous to normal behavior ratio of approximately 3:2.
Figure 3: (a) Overview of our Cross-Modal Pose-aware (CMP) framework, composed of (b) a pose-aware image encoder, a text encoder, and a cross encoder. We apply (c) Identity-based Hard Negative Mining to form challenging negative pairs during training, followed by feature extraction, contrastive learning, and processing by the cross encoder with Image-Text Matching (ITM) and Mask Language Modeling (MLM) heads. The final step computes a cross-modal loss including ITM and MLM components.